### v1

# Revealing Superconducting Chiral Edge Modes via Resistance Distributions

Abstract
Inducing superconducting correlations in quantum anomalous Hall (QAH) states offers a promising route to realize topological superconductivity with chiral Majorana edge modes. However, the definitive identification of these modes is challenging. Here we propose detecting superconducting chiral edge modes via the probability distribution of the resistance, or equivalently the charge transmission of QAH-superconductor heterojunctions. Remarkably, the distribution for coherent edge exhibits distinct characteristics for different topological superconducting phases in sufficiently long junctions, and this difference remains robust against weak decoherence. These findings provide new insights into transport phenomena beyond the clean limit and highlight the resistance distribution as a compelling signature for distinguishing topological superconducting phases.
The superconducting chiral edge modes, which are unidirectional and dissipationless Bogoliubov edge states, could emerge along the boundaries of a two-dimensional chiral topological superconductor (TSC) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] and have potential applications in quantum information [11, 12, 13, 14, 15, 16]. A promising approach to realizing these modes is to induce superconducting pair correlation in the chiral edge states of quantum Hall (QH) or quantum anomalous Hall (QAH) insulators via the proximity effect [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Distinctive transport phenomena have been observed in the QH-SC and QAH-SC junctions [24, 25, 27, 26, 28] due to the charge-neutral nature of these edge modes [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. A significant goal is to achieve $p+ip$ TSC, which hosts a single chiral Majorana edge mode exhibiting quantized transport and interference signatures [34, 35, 36, 37, 52]. However, its experimental confirmation remains elusive. At the same time, disorder and decoherence are inherent in realistic experimental devices [28, 26], making it essential to study transport signatures beyond the idealized clean limit. In particular, understanding the resistance distribution under these effects is vital [43, 48, 49], especially as such distributions can now be measured directly in experiments [26]. Theoretically, an intriguing question is whether these distributions reveal novel behaviors due to the presence of exotic Majorana edge states. Experimentally, it is crucial to determine whether the distribution, beyond just its mean value, provides stronger evidence to distinguish between different topological phases in SC.
In this Letter, we present a comprehensive analytical investigation of disorder-induced resistance distributions in QAH-SC junctions, supported by extensive numerical simulations. Using the transfer matrix formalism, we show that the quantum transport of edge modes under disorder can be intuitively described as random rotations on the Bloch sphere. This framework offers a more rigorous approach than that of [43], which relied on simplifications from renormalization group analysis. We specifically consider QAH-SC heterojunctions commonly implemented in experiments as illustrated in Fig. 1, where the QAH insulator has a Chern number $C=1$, hosting a single chiral electron edge mode [53, 54, 55, 56, 57], while the SC region is characterized by a Bogoliubov-de Gennes (BdG) Chern number $N=1$ or $N=2$, corresponding to transport mediated by one or two chiral Bogoliubov edge modes (CBEM), respectively. The key results are summarized in Table 1, which outlines the probability distributions of charge transmission (defined in Eq. (1)) across different scenarios. Notably, for sufficiently long junctions, the distributions for $N=1$ and $N=2$ CBEM exhibit distinct characteristics. Moreover, these differences remain robust against decoherence, suggesting that resistance distributions serve as a compelling signature to distinguish between different topological phases of SC.
Table 1: Summary of the probability distributions of charge transmission fraction $T$ in two commonly implemented heterojunction configurations. $f(T)$: the probability density function of $T$. $\mathcal{N}$: normal distribution. $\chi^{2}$: chi-squared distribution. $U[-1,1]$: uniform distribution over $[-1,1]$.
Configuration CBEM Length Cases $f(T)$
QAH-SC-QAH $N=2$ short (i) $\mathcal{N}$ or $\chi^{2}$
long (ii) $U[-1,1]$
$N=1$ (iii) $\mathcal{N}$
QAH-SC $N=2$ short (i) $\mathcal{N}$ or $\chi^{2}$
long (ii) $U[-1,1]$
$N=1$ short (i) $\mathcal{N}$ or $\chi^{2}$
long (iv) generalized arcsine
Edge transport with $N=2$ CBEM. By stacking a normal SC onto the middle region of a $C=1$ QAH insulator, one can experimentally implement a QAH-SC-QAH heterojunction in Fig. 1(a). We first consider the middle region of the heterojunction to be in the same topological phase as the $C=1$ QAH state with broken charge $U(1)$ symmetry, namely a TSC with BdG Chern number $N=2$, which would be the case when the SC proximity gap is smaller than the insulating QAH bulk gap [6, 37]. In Nambu basis $(c_{k}^{\dagger},~{}c_{-k})^{\text{T}}$, the effective edge Hamiltonian for TSC is described as $H_{\text{edge}}=vk\tau_{0}-\mu\tau_{z}+\Delta\tau_{x}$, where $v$ is the edge mode velocity, $\mu$ is the chemical potential, and $\Delta=\Delta_{0}k$ is the SC pairing amplitude. $\tau_{x,y,z}$ are Pauli matrices in Nambu space. Accordingly, the single chiral electron edge mode of the QAH state splits into a particle-hole conjugate pair of CBEM of the TSC state, also knows as chiral Andreev edge modes [24]. In Fig. 1(a), normal transmission and Andreev transmission [58] refer to the processes where an electron from the QAH region enters the SC and exits as either an electron or a hole, with transmission coefficients denoted by $T_{\text{ee}}$ and $T_{\text{eh}}$, respectively. We define the charge transmission fraction [59] from lead 2 to 3 as,
$T\equiv T_{\text{ee}}-T_{\text{eh}}.$ (1)
The presence of static disorder induces spatial variations in the potential and pairing amplitude experienced by the edge modes. This effect is effectively described by random fluctuations of the parameters $(v,\mu,\Delta_{0})$ in $H_{\text{edge}}$, making $T$ a random variable governed by a probability distribution—this forms the central focus of our study. Experimentally, this distribution can be observed by varying the magnetic field or the electron density [60, 48, 26]. We divide the SC region into $L$ segments labeled by the index $\ell=1,2,...,L$, where $L\propto\mathcal{L}/\xi_{d}$. Here $\mathcal{L}$ denotes the junction length and $\xi_{d}$ is the disorder correlation length. Within each segment, the parameters $(v,\mu,\Delta_{0})$ remain constant in the local Hamiltonian $H_{\ell}$, as depicted in Fig. 1(b). The final state is obtained via the transfer matrix method [33, 61, 62]
$\Psi_{f}=(\psi_{\text{e}},\psi_{\text{h}})^{\text{T}}=\mathbb{T}_{L}\cdots\mathbb{T}_{2}\mathbb{T}_{1}\Psi_{i}\equiv\mathbb{T}\Psi_{i},$ (2)
where $\Psi_{i}=(1,0)^{\text{T}}$ is the incident state. The transfer matrix for the $\ell$-th SC segment $\mathbb{T}_{\ell}$ is given by $\mathbb{T}_{\ell}=P_{\ell}^{-1}\Lambda_{\ell}P_{\ell}$, where $\Lambda_{\ell}$ represents the propagation matrix of the segment and $P_{\ell}$ transforms the Nambu basis into the eigenbasis of $H_{\ell}$. $\mathbb{T}_{\ell}$ is intuitively interpreted as a rotation on the Bloch sphere, with a rotation angle $\alpha_{\ell}$ around an axis $\mathbf{n}_{\ell}$, as depicted in Fig. 1(c). Here we consider $L>2\ell_{d}/a$, where $\ell_{d}$ is the decay length of CBEM wavefuntion, and $a$ is the lattice constant. Therefore, the crossed Andreev reflection and elastic co-tunneling processes are neglected [63, 46, 64, 65]. As detailed in [62], both $\alpha_{\ell}$ and $\mathbf{n}_{\ell}$ exhibit small fluctuations, with standard deviations satisfying $\delta\alpha_{\ell}\ll 2\pi$ and $|\delta\mathbf{n}_{\ell}|\ll 1$. The cumulative effect of these random rotations can be approximated as a single rotation by an angle $\alpha$ around an axis $\mathbf{n}\equiv(n_{x},n_{y},n_{z})$, such that $\mathbb{T}\propto e^{i\alpha(\mathbf{n}\cdot\bm{\sigma})/2}$, where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are Pauli matrices. The final state then takes the form $\Psi_{f}=(\cos{(\alpha/2)}+in_{z}\sin{(\alpha/2)},i(n_{x}+in_{y})\sin{(\alpha/2)})^{\text{T}}$, and the charge transmission fraction is given by $T=|\psi_{\text{e}}|^{2}-|\psi_{\text{h}}|^{2}=n_{z}^{2}+(1-n_{z}^{2})\cos{\alpha}$.
Figure 1: (a) The QAH-SC-QAH junction with $N=2$ CBEM. The orange and dashed blue arrows represent the chiral electron edge mode of QAH state and the CBEM of TSC state, respectively. (b) Transfer matrix construction for the junction shown in (a). (c) Evolution of the state on the Bloch sphere. Both the axis and rotation angle exhibit weak fluctuations. The north (south) pole on the Bloch sphere represents a pure electron (hole) state. (d,e) The QAH-SC-QAH junction with $N=1$ CBEM, and the corresponding transfer matrix construction. (f) The QAH-SC junction with $N=1$ CBEM.
The distribution of $T$ exhibits distinct yet general behaviors depending on the length of SC region, as analyzed in the following cases.
Case (i): For a relatively short QAH-SC-QAH junction (small $L$ as $2\ell_{d}/a<L<10^{2}$) in Fig. 1(a), $T$ approximately follows either a normal distribution ($\mathcal{N}$), or a chi-squared distribution ($\chi^{2}$).
In this regime, fluctuations in the rotation axis are negligible, allowing us to approximate $\mathbf{n}\approx\overline{\mathbf{n}}_{\ell}$. As a result, the state undergoes sequential rotations around a nearly fixed axis, with each segment contributing a random rotation angle. The total rotation angle is $\alpha=\sum_{\ell=1}^{L}\alpha_{\ell}$. By the central limit theorem, after multiple segments, $\alpha$ follows a normal distribution with mean $\overline{\alpha}=L\overline{\alpha}_{\ell}$ and standard deviation $\delta{\alpha}=\sqrt{L}\delta{\alpha}_{\ell}$, irrespective of the specific distribution of the individual rotation angles $\alpha_{\ell}$. When $L$, and thus $\delta\alpha$, is small, we can expand $\alpha$ around $\overline{\alpha}$ up to second order: $\cos{\alpha}\approx\cos{\overline{\alpha}}-\sin{\overline{\alpha}}\Delta\alpha-\cos{\overline{\alpha}}(\Delta\alpha)^{2}/2$, where $\Delta\alpha$ follows a normal distribution. Therefore, $T$ approximately follows a normal distribution when $\sin{\overline{\alpha}}$ is not close to zero, and a chi-squared distribution when $\sin{\overline{\alpha}}\approx 0$ [62]. As the junction length varies, the distribution of $T$ oscillates between normal and chi-squared distributions. This behavior extends beyond the results in [48], where only a normal distribution was proposed.
Case (ii): For a long QAH-SC-QAH junction (large $L$ as $L>10^{4}$) in Fig. 1(a), $T$ approximates a continous uniform distribution over $[-1,1]$, denoted as $U[-1,1]$.
As $L$ increases, the fluctuations of the rotation axis become significant, causing the state to spread across the entire Bloch sphere, as illustrated in Fig. 1(c). The ergodicity theorem [66] states that a point undergoing repeated rotations about two noncollinear axes eventually becomes uniformly distributed over the sphere. Since the fluctuating rotation axes in our system satisfy this condition, the final state $\Psi_{f}$ becomes equidistributed on the Bloch sphere, leading to $T\sim U[-1,1]$. A more rigorous derivation is provided in [62].
Edge transport with $N=1$ CBEM. Next, we consider the case where the middle SC region of the heterojunction has a BdG Chern number $N=1$, hosting a single self-conjugate CBEM, also known as chiral Majorana edge mode [6, 37]. As shown in Fig. 1(d), two incoherent incident charged modes, originating from different sources and labeled $a$ and $b$, decompose into four Majorana edge modes, labeled $\gamma_{1}$ through $\gamma_{4}$. Among them, $\gamma_{2}$ and $\gamma_{3}$ undergo braiding, after which all four modes recombine into two outgoing charged modes, labeled $c$ and $d$. This process, depicted in Fig. 1(e), is described by
$\Psi_{f}=Q^{\prime}M^{\prime}S_{23}\Lambda MQ\Psi_{i}.$ (3)
Here, $\Psi_{i}$ is a four-dimensional vector representing the initial state. The transformation sequence, proceeding from right to left, is as follows: $Q$ accounts for propagation near the corners upon entering the SC region, $M$ transforms the Nambu basis into the Majorana basis, $\Lambda$ describes propagation within the SC region in the Majorana basis, $S_{23}$ captures the braiding of $\gamma_{2}$ and $\gamma_{3}$, $M^{\prime}$ rotates the Majorana basis back to the Nambu basis, and $Q^{\prime}$ describes propagation near the corners before exiting the SC region. The matrix $\Lambda$ is diagonal, with random phase factors reflecting the well-separated and decoupled nature of the Majorana modes away from the SC region’s corners. However, near these corners, the Majorana modes overlap, leading to non-trivial coupling. This coupling takes the form $\eta\tau_{z}$ in the Nambu basis, where $\eta$ is a disorder-dependent coupling coefficient [40, 43]. This coupling term corresponds to the chemical potential term in $H_{\text{edge}}$. Consequently, the effect of random coupling is captured by $Q=\text{diag}(\mathbb{T}^{a},\mathbb{T}^{b})$ and $Q^{\prime}=\text{diag}(\mathbb{T}^{c},\mathbb{T}^{d})$, where $\mathbb{T}^{\zeta}$ represents the total transfer matrix at corner $\zeta=a,b,c,d$, structured similarly to the transfer matrix product in Eq. (2).
Case (iii): For the QAH-SC-QAH junction with $N=1$ CBEM shown in Fig. 1(d), $T$ consistently follows a normal distribution, regardless of the junction length.
The normal distribution of $T$ stems from two key factors. First, the incoherent nature of the incident electrons ensures that the phase factors in $\Lambda$ do not affect $T$, making it independent of the junction length. Second, coupling occurs only near the corners, leading to a small effective $L$ when computing $\mathbb{T}^{\zeta}$. As a result, each $\mathbb{T}^{\zeta}$ can be treated as a random rotation around a fixed axis, analogous to Case (i). A detailed derivation of this result is provided in [62].
Then we examine an alternative junction configuration [24, 25, 27, 26, 28], depicted in Fig. 1(f). This setup, known as a QAH-SC junction, is realized by stacking a normal SC onto the edge region of a $C=1$ QAH insulator. When the SC region has a BdG Chern number $N=2$, the charge transmission fraction $T$ from lead 2 to 3 in QAH-SC junction matches that of the QAH-SC-QAH junction in Fig. 1(a). Therefore, we focus on the SC region has $N=1$. In this configuration, the incident charged mode $a$ splits into two Majorana modes, $\gamma_{1}$ and $\gamma_{2}$, within the SC region, which subsequently recombine and emerge as the outgoing charged mode $b$. A key distinction between this setup and the QAH-SC-QAH junction in Fig. 1(d) is that in the QAH-SC junction, the Majorana modes originate from the same source, preserving their coherence. This coherence leads to different transport behaviors of $T$, depending on the junction length. Specifically, the final state can be expressed as
$\Psi_{f}=\mathbb{T}^{b}M^{\prime}\Lambda M\mathbb{T}^{a}\Psi_{i}.$ (4)
Here, fluctuations near the corners are negligible compared to those along the edges of the SC region. This allows us to approximate the total transfer matrix as $\overline{\mathbb{T}}^{-1}M^{\prime}\Lambda M\overline{\mathbb{T}}$, where $\overline{\mathbb{T}}$ represents the average transfer matrix at the corner. As a result, the entire process can be interpreted as a random rotation by an angle $\alpha$ around a fixed axis $\mathbf{n}$, with $\alpha$ and $\mathbf{n}$ determined by $\Lambda$ and $M\overline{\mathbb{T}}$, respectively. For small $L$, the fluctuation of $\alpha$ remains small, making the system behave similarly to Case (i). However, as $L$ increases, the system exhibits qualitatively different behavior.
Case (iv): For a long QAH-SC junction (large $L$ as $L>10^{4}$) with $N=1$ CBEM in Fig. 1(f), $T$ approximately follows a generalized arcsine distribution.
As discussed in Case (i), $\delta\alpha\propto\sqrt{L}$, which becomes significant for large $L$. Therefore, $\cos{\alpha}$ will converge to the standard arcsine distribution [67], then $T$ obeys a shifted and rescaled variant of the standard arcsine distribution, with the probability density function
$f(T)=\begin{cases}\frac{1}{\pi\sqrt{(1-A)^{2}-(T-A)^{2}}}\ ,&(T\in[1-2A,1])\\ 0\ ,&(T\notin[1-2A,1])\end{cases}$ (5)
where $A=n_{z}^{2}$. This distribution is characterized by a U-shaped profile that peaks at the maximal value 1 and the minimal value ($1-2A$) of $T$.
Finally, when the SC region of the junction is topologically trivial ($N=0$), two CBEM emerge along the interface between the SC and QAH. Therefore, the edge mode transport characteristics closely resemble those of the $N=2$ case [62].
Numerical calculations. To validate the analytical results presented above, we numerically compute $T$ in each case. We implemented two-dimensional tight-binding Hamiltonians of both QAH and SC for the configurations in Fig. 1 (a,d,f) on a rectangular geometry. The disorder is introduced through a uniformly distributed onsite potential with $\xi_{d}=a\equiv 1$. The charge transmission $T$ is then calculated using the recursive Green’s function method [68, 69, 70, 71, 72, 73]. Additional numerical details are provided in [62].
Figure 2: Mean and distribution of the charge transmission fraction $T$ for the junction in Fig. 1(a). (a,b) $T$ vs $L$, where solid lines and shaded region indicate the mean and standard deviation, respectively. (c-e) The distributions of $T$ for different $L$, with histograms representing numerical results and red lines showing fitted distributions. (c) The distribution is fitted with a normal distribution $f(T)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(T-E)^{2}}{2\sigma^{2}}}$. (d) The distribution is fitted with a chi-squared distribution $f(T)=\frac{1}{\sigma\sqrt{\pi(1-T)}}e^{-\frac{1-T}{\sigma^{2}}}$. (e) The mean and standard deviation of $T$ are $8.54\times 10^{-3}$ and 0.582, respectively, which match those of a uniform distribution $U[-1,1]$.
The results for the junction configuration of Fig. 1(a) are presented in Fig. 2. Here, $\overline{T}$ denotes the mean charge transmission fraction from lead 2 to 3. When disorder is introduced into the SC region, $\overline{T}$ exhibits oscillations with a decaying amplitude as the junction length $L$ increases, eventually converging to zero. Meanwhile, the fluctuations in $T$ increase, as shown in Fig. 2(a,b). For $L\gtrsim 2\ell_{d}$, $T$ follows a normal or chi-squared distribution, as illustrated in Fig. 2(c) and Fig. 2(d), respectively, in accordance with Case (i). For sufficiently large $L$, such that $L\gtrsim 10^{4}$, the distribution of $T$ approaches a continuous uniform distribution $U[-1,1]$, as shown in Fig. 2(e), corresponding to Case (ii).
Figure 3: (a,b) Numerical results of $T$ for junction in Fig. 1(d). (a) $\overline{T}$ vs $L$, where $\varepsilon$ denotes the incident electron energy. The solid lines and shading indicate the mean and standard deviation, respectively. (b) Distribution of $T$ for $\varepsilon=0.28$, which follows a normal distribution depicted as red line. (c,d) Numerical results for junction of Fig. 1(f). (c) $\overline{T}$ vs $L$. (d) Distribution of $T$, with the red line showing the fitting distribution as a generalized arcsine distribution in Eq. (5).
For the junction configuration of Fig. 1(d), the results are presented in Fig. 3(a,b). When $L\gtrsim 2\ell_{d}$, both the mean and fluctuations of $T$ remain constant as $L$ varies, with $\overline{T}$ aligning with its value in the clean limit. Furthermore, the distribution of $T$ follows a normal distribution, consistent with Case (iii), confirming that this distribution arises from the random coupling of Majorana modes near the junction corners. Next, the results for the QAH-SC junction of Fig. 1(f) are shown in Fig. 3(c,d). While $\overline{T}$ exhibits oscillations with decaying amplitude—similar to the $N=2$ case—it does not necessarily converge to zero, depending on the incident electron energy $\varepsilon$. Furthermore, for sufficiently large $L$, such that $L\gtrsim 10^{4}$, the distribution transitions to a generalized arcsine form, as predicted in Case (iv). The left peak is broadened due to the neglected fluctuations of the rotation axis in the analytical analysis, which introduce slight variations in the lower bound of the distribution.
Decoherence. Two additional factors, loss and decoherence, play a significant role in shaping experimental observations beyond the effects of disorder [26]. Loss occurs when particles tunnel into magnetic vortices and escape through the vortex line to the substrate [48, 46, 74]. This process reduces the norm of $\Psi_{f}$, thus compressing the distribution of $T$ without altering its overall shape. Decoherence, on the other hand, results from inelastic scattering with other gapless degrees of freedom, which can be induced by magnetic vortices, ineffective SC proximity, or thermal fluctuations [49]. When the decoherence effect is weak, the effect of decoherence can be conceptualized as the edge mode that has a probability $p$ of undergoing a single inelastic scattering event during its propagation. The resulting effective charge transmission fraction for different scenarios in Fig. 1(a,d,f) is shown in Fig. 4, where the distribution for $N=2$ CBEM remains distinct from that for $N=1$ CBEM, indicating that the characteristic behavior of the distribution in the coherent process persists under sufficiently weak decoherence. As the decoherence effect intensifies, characterized by an increase in $p$ and the occurrence of multiple scattering events, the distribution for the scenario in Fig. 1(a) will become logarithmic divergent peaked at $T=0$, while the scenarios in Fig. 1(d,f) become bell-shaped peaked at finite $T$ [62].
Figure 4: (a,b,c) Decoherence effect in scenarios Fig. 1(a,d,f), respectively. The distribution is calculated from Monte Carlo sampling [62], and we set $p=0.4$. We set the distribution of $T$ without decoherence follows $U[-1,1]$, $\mathcal{N}(0.3,0.05^{2})$ and arcsin distribution with $A=0.6$ for Fig. 1(a,d,f), respectively.
Discussions. The experimentally measurable resistances are determined by $T$. Taking QAH-SC junction shown in Fig. 1(f) as an example. The current is from lead 1 to the grounded SC, namely $I_{1}=-I_{s}=I$, and voltage leads are on electrodes 2, 3, 4, and 5. The downstream resistance [19, 25, 28] is given by the generalized Landauer-Büttiker formula [75]
$R_{d}\equiv R_{3s}=\frac{V_{3}-V_{s}}{I}=\frac{h}{e^{2}}\frac{T}{1-T},$ (6)
where $h$ is the Planck constant, $e$ is the electrical charge. Thus, $T$ can be directly extracted from the measured $R_{d}$. Other nonlocal resistances in different scenarios are provided in [62].
Our theory shows that the probability distribution of charge transmission $T$ provides a signature distinguish between the different topological phases of the SC region. As summarized in Table 1, for sufficiently long junctions, the distributions for $N=1$ and $N=2$ CBEM exhibit distinct characteristics. Moreover, these differences remain robust against weak decoherence. Given the stark differences between these distributions, they provide stronger evidence for phase differentiation, offering extra insight than mean alone, which may be insufficient to distinguish between topological phases of SC region [28].
Acknowledgements.
Acknowledgment. We thank Yuanbo Zhang for enlightening discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01.